Jason Thomas has saved $3,500 towards a down payment on a new truck. He wants to have a total of $20,000 and plans to buy in 3 years. How much will he have to save at the end of each month if he can earn 2.5% on his savings?

Multiple Choice

  • $433.64

  • $635.25

  • $636.58

  • $434.55
Answer #1

$434.55

Step-1:Future value of current saving
Future Value = P*(1+i)^n Where,
= 3500*(1+0.0020833)^36 P = $ 3,500.00
= $   3,772.30 i = 2.5%/12 = 0.0020833
n = 3*12 = 36
Step-2:Required monthly saving
Monthly saving = Future value of monthly saving / Future value of annuity of 1
= $ 16,227.70 / 37.34403
= $       434.55
Working:
# 1
Total amount required $ 20,000.00
Less future value of current saving $   3,772.30
Balance required $ 16,227.70
# 2
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= 37.3440294 i = 2.5%/12 = 0.0020833
n = 3*12 = 36